# GR-MG: Leveraging Partially Annotated Data via Multi-Modal Goal Conditioned Policy

### v1

@@ Abstract
The robotics community has consistently aimed to achieve generalizable robot manipulation with flexible natural language instructions. One of the primary challenges is that obtaining robot data fully annotated with both actions and texts is time-consuming and labor-intensive. However, partially annotated data, such as human activity videos without action labels and robot play data without language labels, is much easier to collect. Can we leverage these data to enhance the generalization capability of robots? In this paper, we propose GR-MG, a novel method which supports conditioning on both a language instruction and a goal image. During training, GR-MG samples goal images from trajectories and conditions on both the text and the goal image or solely on the image when text is unavailable. During inference, where only the text is provided, GR-MG generates the goal image via a diffusion-based image-editing model and condition on both the text and the generated image. This approach enables GR-MG to leverage large amounts of partially annotated data while still using language to flexibly specify tasks. To generate accurate goal images, we propose a novel progress-guided goal image generation model which injects task progress information into the generation process, significantly improving the fidelity and the performance. In simulation experiments, GR-MG improves the average number of tasks completed in a row of 5 from 3.35 to 4.04. In real-robot experiments, GR-MG is able to perform 47 different tasks and improves the success rate from 62.5% to 75.0% and 42.4% to 57.6% in simple and generalization settings, respectively. Code and checkpoints will be available at the project page: https://gr-mg.github.io/.
Figure 1: Overview. GR-MG consists of two modules: the progress-guided goal image generation model and the multi-modal goal-conditioned policy. The former generates a goal image based on the current observation, language instruction, and task progress. The latter predicts task progress and actions based on the goal image produced by the former. Data without action labels can be used to train the goal image generation model, while the multi-modal goal-conditioned policy can be trained on robot play data that lacks language annotations. Fully annotated data will be used to train both modules.
@@ I INTRODUCTION
The robotics research community is striving to achieve generalizable robot manipulation using language-based instructions. Among various methods, imitation learning from human demonstrations is one of the most promising endeavors[1, 2, 3, 4, 5]. However, human demonstration data are scarce. And the process of collecting human demonstrations with action and language labels is time-consuming and labor-intensive. On the other hand, robot play data without language labeling is a scalable source of data. It requires no video clipping or language labeling and can be collected without human constant supervision [6]. These data can be automatically labeled with hindsight goal relabeling [7]: any frame in a video can be used as a goal image to condition the policy for predicting actions to evolve towards this frame from previous frames. Furthermore, there is a substantial collection of text-annotated human activity videos without action labels on the internet. These data contain valuable information about how the agent should move to change the environment according to the language description. Can we develop a policy to effectively leverage all the above partially annotated data?
@@
Previous methods have ventured into this domain, but most were limited to using data that lacked either language labels or action labels [4, 7, 8, 9, 10]. Recent initiatives introduce diffusion models for generating goal images [11] or future videos [12, 13]. The generated image or video is then used as inputs for goal-conditioned policies or inverse dynamics models to predict actions, enabling the use of all the above-mentioned partially annotated data.
However, these approaches often overlook crucial information, such as task progress, during the goal generation phase. This omission can lead to inaccurate generated goals which significantly affects the subsequent action predictions. Furthermore, these policies rely solely on images or videos, making them brittle in the case where the generated goal is inaccurate.
@@
To tackle these issues, we introduce GR-MG, a model designed to support multi-modal goals. It comprises two modules: a progress-guided goal image generation model and a multi-modal goal-conditioned policy. Datasets without action labels (e.g., text-annotated internet videos) can be used to train the goal image generation model. Datasets without language labels (e.g., robot play data) can be used to train the policy, with a null string as the language condition. Given that robot manipulation is a sequential decision-making process, where the task progress information can enhance prediction accuracy, we incorporate a novel task progress condition into our goal image generation model. This significantly improves the performance of both goal image generation and thus action prediction. During the training of the policy, we sample goal images from trajectories and condition the policy on both the goal image and the text instruction or solely on the goal image if text is unavailable. During inference, the policy utilizes both the language instruction and the goal image generated from the goal image generation model. Since GR-MG is conditioned on both the text and the goal image, the policy can still rely on the language condition to guide action prediction even if the generated goal image is inaccurate, substantially improving the robustness of the model.
@@
We perform extensive experiments in the challenging CALVIN simulation benchmark [14] and a real-robot platform. In CALVIN, without using additional partially annotated data, GR-MG significantly outperforms all the comparing state-of-the-art methods in a zero-shot generalization setting, improving the success rate of completing 1 and 5 tasks in a row from 93.8% to 96.8% and 41.2% to 64.4%, respectively. When we add robot play data into training, GR-MG achieves average length of 3.11 with only 10% of provided fully annotated data, which is even better than the competitive GR-1 [4] baseline using all the data. In real-robot experiments, we evaluate GR-MG in both simple and 3 challenging generalization settings. In total, GR-MG can perform 47 tasks and improves the average success rate from 62.5% to 75.0% in the simple setting and from 42.4% to 57.6% in generalization settings. In summary, the contribution of this paper is threefold:
@@
  • We propose a novel Generative Robot Policy with Multi-modal Goals (GR-MG), which is able to leverage both text-annotated videos and robot play data during training.
@@
  • We introduce the task progress condition in goal image generation, which substantially improves the accuracy of the generated goal image.
@@
  • We perform extensive experiments and ablation studies in simulation and the real world to verify the effectiveness of GR-MG in both simple and challenging generalization settings.
@@
Figure 2: Network Architecture. We use a diffusion model to generate the goal image based on the current image, language instruction, and task progress. The generated goal image is then sent to the multi-modal goal-conditioned policy, which takes as inputs the language instruction, sequences of visual observations, and robot states. The policy is a GPT-style transformer. We insert query tokens [PROG], [OBS], and [ACT] after the input tokens to predict actions, future images, and the progress of the current state, respectively. The progress is fed back to the goal generation model. In this figure, we show the setup with two camera views: a static view and a view captured from a camera mounted on the end-effector.
@@ II-A Leveraging Multi-Modal Goals in Policy Learning
Language is probably the most flexible and intuitive way for a human to specify tasks for a robot [1, 2, 4, 5, 15]. However, language and visual signals exist in different domains. And the information contained in a language instruction may be too abstract for a visual policy to comprehend. A line of research explores goal-image conditioned policies [16]. But obtaining goal images during rollouts can be challenging. Recently, some studies have proposed leveraging both language and goal images as conditions during training. They utilize data with language labels alongside robot play data that without language labels [7, 17, 18, 3]. A common approach is to project both the goal image and the language into a shared latent space to condition the policy. However, only the language is used as the condition during inference. GR-MG differs from these methods in that it uses both languages and goal images during inference and training. The goal image is generated via a diffusion model from the language and current observation. The most related work to GR-MG is Susie [11]. It also uses a diffusion model to generate the goal image. The goal image is then fed into a diffusion policy [19] for action prediction. GR-MG follows a similar paradigm but differs in two key aspects. First, it introduces a novel task progress condition in goal image generation, significantly enhancing the accuracy of the generated goal image. Second, GR-MG conditions the policy using both the language and the goal image, whereas Susie relies solely on the generated goal images. This design enables GR-MG to address situations where the generated goal image is inaccurate, thereby improving performance in challenging out-of-distribution environments.
II-B Leveraging Partially Annotated Data in Policy Learning
@@
Given that collecting fully annotated robot data with action and language labels is time-consuming, many studies turn to alternative sources of data with partial annotations. One approach involves learning useful representations from text-annotated internet videos, which are subsequently used for downstream policy learning [20, 21]. Recently, RT-2 [2] introduced a vision-language-action model that incorporates both robot data and large-scale vision-language data during training, demonstrating powerful generalization capabilities. Another popular method is to learn a video model or world model from internet video data to predict future videos [22, 13, 12, 8]. An inverse dynamics model or goal-conditioned policy is then employed to predict actions based on the predicted video. Recently, Wu et al. [4] proposed GR-1, which is pre-trained on large-scale text-annotated videos from the internet and then fine-tuned on robot trajectories with language labels. GR-MG differs from GR-1 in two aspects. First, GR-MG leverages robot play data during training, while GR-1 does not. Second, GR-MG utilizes both the language and the goal image as conditions, whereas GR-1 relies solely on the language. These differences significantly enhance GR-MG’s generalization capabilities, as we will demonstrate in Sec. IV.
@@ III Method
We aim to learn a policy $\pi$ to address the problem of language-conditioned visual robot manipulation. Specifically, the policy takes as inputs the following. 1) A language instruction $l$ that describes the task to be accomplished. 2) Observation images $\mathbf{o}_{t-h:t}$ which include RGB images from the current and previous $h$ timesteps. A static global view camera and a hand-mounted camera are used to capture these images. 3) Robot states $\mathbf{s}_{t-h:t}$ which include the 6-Dof pose of the end-effector and the binary gripper state of the current and previous $h$ steps. The policy outputs an action trajectory $\mathbf{a}_{t}$ in an end-to-end manner:
@@ III-A Data
The training of a language-conditioned policy requires fully annotated robot trajectories $\tau$ which contain both action/state labels and language labels:
@@
However, collecting fully annotated data is time-consuming. The motivation behind GR-MG is to utilize partially annotated data. One such data includes videos without action labels, which can be collected at scale from the internet and various public datasets [23, 24, 25]. Another type of partially annotated data consists of robot play data without language labels. These can be gathered on a large scale by allowing robots to execute a policy without constant human supervision [6].
@@ III-B Network Architecture
The network architecture of GR-MG is illustrated in Fig. 2. It consists of two modules: a progress-guided goal image generation model and a multi-modal goal conditioned policy.
@@ III-B1 Progress-guided Goal Image Generation Model
This module generates the goal image based on the current observation image, a language description of the task, and the task progress. Specifically, we use InstructPix2Pix [26], a diffusion-based image-editing model, as the network. Following the approach in Susie [11], we generate a sub-goal representing an intermediate state rather than the final state, updating this sub-goal at fixed time intervals during the rollout. However, unlike Susie which generates the goal image based solely on the current image and the language instruction, we incorporate task progress information in the image generation process. The insight is that robot manipulation is a sequential decision process instead of static image editing. Additionally, a current image can be ambiguous in the case where there are identical observation throughout the rollout of a task, e.g., moving an object back-and-forth. The progress information can provide valuable global information for the goal image generation model. During training, the progress of a frame can be easily obtained from its timestep in the trajectory (or video). During inference, such information is not accessible. To address this, we train the policy to predict task progress alongside the action trajectory, as we will outline below. To incorporate progress information into the generation model, we append a sentence to the task description, e.g., "pick up the red block. And 60% of the instruction has been completed.". We use T5-Base [27] to encode the text.
III-B2 Multi-modal Goal Conditioned Policy
@@
Following GR-1 [4], we employ a similar GPT-style transformer network [28]. For completeness, we briefly review GR-1 here. GR-1 takes as input a language instruction, a sequence of observation images, and a sequence of robot states. The model then outputs an action and a prediction of the future image. The language is encoded using a text encoder [29]; images are encoded via MAE [30]; robot states are encoded through linear layers. To predict actions, GR-1 includes a [ACT] query token, and for future image prediction, it learns multiple [OBS] query tokens. The embeddings of different signals and the query tokens are sequenced by timestep and input into the transformer. For more details, we refer readers to [4].
@@
The difference between GR-MG and GR-1 is threefold. Firstly, GR-MG is a multi-modal goal-conditioned policy which takes both a goal image and a language instruction as the condition. GR-1 only takes the language. To inject the goal image condition, we use MAE [30] to tokenize the goal image generated by the goal image generation model and append the tokens at the front of all the input tokens (Fig. 2). The goal image tokens can be attended by all the subsequent tokens for conditioning. Secondly, in order to predict task progress, a [PROG] query token is included as shown in Fig. 2. The output embedding of the token is passed through linear layers to regress the progress value. Finally, we follow recent work [31] to predict action trajectories instead of a single action as in GR-1 via a conditional variational autoencoder (cVAE) [32, 33]. Specifically, we use a VAE encoder to encode the action trajectory into a style vector embedding. We concatenate the style vector embedding, the output embedding of the [ACT] token and $k$ learnable token embeddings together and input them to a transformer to predict an action trajectory of $k$ steps.

### v2

# GR-MG: Leveraging Partially-Annotated Data via Multi-Modal Goal-Conditioned Policy

@@ Abstract
The robotics community has consistently aimed to achieve generalizable robot manipulation with flexible natural language instructions. One primary challenge is that obtaining robot trajectories fully annotated with both actions and texts is time-consuming and labor-intensive. However, partially-annotated data, such as human activity videos without action labels and robot trajectories without text labels, are much easier to collect. Can we leverage these data to enhance the generalization capabilities of robots? In this paper, we propose GR-MG, a novel method which supports conditioning on a text instruction and a goal image. During training, GR-MG samples goal images from trajectories and conditions on both the text and the goal image or solely on the image when text is not available. During inference, where only the text is provided, GR-MG generates the goal image via a diffusion-based image-editing model and conditions on both the text and the generated image. This approach enables GR-MG to leverage large amounts of partially-annotated data while still using languages to flexibly specify tasks. To generate accurate goal images, we propose a novel progress-guided goal image generation model which injects task progress information into the generation process. In simulation experiments, GR-MG improves the average number of tasks completed in a row of 5 from 3.35 to 4.04. In real-robot experiments, GR-MG is able to perform 58 different tasks and improves the success rate from 68.7% to 78.1% and 44.4% to 60.6% in simple and generalization settings, respectively. It also outperforms comparing baseline methods in few-shot learning of novel skills. Video demos, code, and checkpoints are available on the project page: https://gr-mg.github.io/.
Index Terms:
Deep Learning in Grasping and Manipulation, Imitation Learning, Learning from Demonstration
Figure 1: Overview. GR-MG consists of two modules: a progress-guided goal image generation model and a multi-modal goal-conditioned policy. The former generates a goal image based on the current observation, text instruction, and task progress. The latter predicts task progress and actions based on the text and the goal image produced by the former. Data without action labels can be used to train the goal image generation model, while the multi-modal goal-conditioned policy can leverage data without text labels. Fully-annotated data is used for training both modules.
@@ I INTRODUCTION
The robotics research community is striving to achieve generalizable robot manipulation using language-based instructions. Among various methods, imitation learning from human demonstrations is one of the most promising endeavors[1, 2, 3, 4, 5]. However, human demonstration data are scarce. And the process of collecting human demonstrations with action and text labels is time-consuming and labor-intensive. On the other hand, trajectories without language annotations are a scalable source of data. They do not require executing pre-defined tasks or text labeling, and can be collected without human constant supervision [6]. These data can be automatically labeled with hindsight goal relabeling [7]: any frame in a video can be used as a goal image to condition the policy for predicting actions to evolve towards this frame from previous frames. Furthermore, there is a substantial collection of text-annotated human activity videos without action labels from various public datasets [8, 9, 10]. These data contain valuable information about how the agent should move to change the environment according to the text description. Can we develop a policy to effectively leverage all the above partially-annotated data?
@@
Previous methods have ventured into this domain, but most were limited to using data without either text labels or action labels [4, 7, 11, 12, 13]. Recent initiatives introduce diffusion models for generating goal images [14] and future videos [15, 16]. The generated image or video is then used as inputs for goal-conditioned policies or inverse dynamics models to predict actions, enabling the use of all the above-mentioned partially-annotated data. However, these approaches often overlook crucial information, such as task progress, during the goal generation phase. This can lead to inaccurate generated goals which may significantly affect the subsequent action prediction. Furthermore, these policies rely solely on the goal image or video for action prediction, making them brittle in the case where the generated goal is inaccurate.
@@
To tackle these issues, we introduce GR-MG, a model designed to support multi-modal goals. It comprises two modules: a progress-guided goal image generation model and a multi-modal goal-conditioned policy. GR-MG is able to leverage data without action labels (e.g., text-annotated human activity videos) to train the goal image generation model along with the fully-annotated robot trajectories. And data without text labels can be used for training the multi-modal goal-conditioned policy. Given that robot manipulation is a sequential decision-making process, we incorporate a novel task progress condition into the goal image generation model. This significantly improves the performance on both the goal image generation and action prediction. During training, we sample goal images from trajectories and condition the policy on both the goal image and the text instruction or solely the goal image if text is unavailable. During inference, the policy utilizes both the text instruction and the goal image generated by the goal image generation model to predict actions. Since GR-MG is conditioned on both the text and goal image, the policy can still rely on the text condition to guide action prediction even if the generated goal image is inaccurate, substantially improving the robustness of the model.
@@
We perform extensive experiments on the challenging CALVIN simulation benchmark [17] and a real-robot platform. In CALVIN, without using additional partially-annotated data, GR-MG significantly outperforms all the comparing state-of-the-art methods in a zero-shot generalization setting. It improves the success rates of completing 1 and 5 tasks in a row from 93.8% to 96.8% and 41.2% to 64.4%, respectively. When incorporating data without text labels into training, GR-MG achieves an average length of 3.11 with only 10% of the provided fully-annotated data, outperforming the competitive GR-1 [4] baseline which uses all the data. In real-robot experiments, we evaluate GR-MG in a simple and four challenging generalization settings. In total, GR-MG is able to perform 58 tasks and improves the average success rate from 68.7% to 78.1% in the simple setting and from 44.4% to 60.6% in generalization settings. GR-MG also surpasses comparing baseline methods in few-shot learning of novel skills. In summary, the contribution of this paper is threefold:
@@
  • We propose a novel Generative Robot Policy with Multi-modal Goals (GR-MG), which is able to leverage both data without action labels and data without text labels during training.
@@
  • We introduce the task progress condition in goal image generation, substantially improving the accuracy of the generated goal image.
@@
  • We perform extensive experiments and ablation studies in simulation and the real world to verify the effectiveness of GR-MG in simple, generalization, and few-shot learning settings.
@@
Figure 2: Network Architecture. We use a diffusion model to generate the goal image based on the current image, text instruction, and task progress. The generated goal image is then sent to the multi-modal goal-conditioned policy, which takes as inputs the text instruction, sequences of observations images and robot states. The policy is a GPT-style transformer. We insert query tokens [PROG], [OBS], and [ACT] after the input tokens to predict actions, future images, and the task progress, respectively. The progress is fed back to the goal image generation model. In this figure, we show the setup with two camera views: a static view and a view captured from a wrist-mounted camera.
@@ II-A Leveraging Multi-Modal Goals in Policy Learning
Language is probably the most flexible and intuitive way for a human to specify tasks for a robot [1, 2, 4, 5, 18]. However, language and visual signals exist in different domains. And the information contained in a language instruction may be abstract for a visual policy to comprehend. A line of research explores goal-image conditioned policies [19, 20]. But obtaining goal images during rollouts is challenging in robot manipulation tasks. Recently, some studies have proposed leveraging both language and visual goals (goal images or videos) as conditions during training [21, 7, 22, 23, 24, 3]. And Mutex [25] is able to leverage multiple modalities for task specification, including speech, text, images, and videos. However, in language-conditioned manipulation, where language is the only condition provided, these methods do not generate goal images. Instead, they rely solely on the language to condition the prediction of actions during inference. GR-MG differs from these methods in that it uses both the language and the goal image during training and inference. And the goal image is generated via a diffusion model from the language and current observation. The most related work to GR-MG is Susie [14]. It also uses a diffusion model to generate the goal image. The goal image is then fed into a diffusion policy [26] for action prediction. GR-MG follows a similar paradigm but differs in two key aspects. First, it introduces a novel task progress condition in goal image generation, significantly enhancing the accuracy of the generated goal image. Second, GR-MG conditions the policy with both the language and goal image, whereas Susie relies solely on the generated goal images. This design enables GR-MG to address situations where the generated goal image is inaccurate, thereby improving its robustness in challenging out-of-distribution settings.
II-B Leveraging Partially-Annotated Data in Policy Learning
@@
Given that collecting fully-annotated robot data with action and language labels is time-consuming, many studies turn to alternative sources of data with partial annotations. One approach involves learning useful representations from text-annotated videos without action labels, which are subsequently used for downstream policy learning [27, 28, 29]. Recently, RT-2 [2] introduces a vision-language-action model that incorporates both robot data and large-scale vision-language data during training, demonstrating powerful generalization capabilities. Another popular method is to learn a video model or world model from existing video data to predict future videos [30, 16, 15, 11]. An inverse dynamics model or goal-conditioned policy is then employed to predict actions based on the predicted video. Recently, Wu et al. [4] proposes GR-1, which is pretrained on large-scale text-annotated videos and then finetuned on robot trajectories with text labels. GR-MG differs from GR-1 in two aspects. First, GR-MG leverages data without text labels during training, while GR-1 does not. Second, GR-MG utilizes both the language and goal image as conditions, whereas GR-1 relies solely on the language. These differences significantly enhance GR-MG’s generalization capabilities.
@@ III Method
We aim to learn a policy $\pi$ for language-conditioned visual robot manipulation. Specifically, the policy takes as inputs a text instruction $l$, a sequence of observation images $\mathbf{o}_{t-h:t}$, and a sequence of robot states $\mathbf{s}_{t-h:t}$. The observation images $\mathbf{o}_{t-h:t}$ include RGB images captured from timestep $t-h$ to $t$. In this paper, we use a static camera and a wrist-mounted camera to capture these images. The robot states $\mathbf{s}_{t-h:t}$ include the 6-DoF poses of the end-effector and the binary gripper states from timestep $t-h$ to $t$. The policy outputs an action trajectory $\mathbf{a}_{t}$ in an end-to-end manner:
@@ III-A Data
Training a language-conditioned policy typically requires fully-annotated robot trajectories $\tau$ which contain both action and text labels:
@@
However, collecting fully-annotated data is time-consuming. The motivation behind GR-MG is to leverage partially-annotated data in training. There are two main types of partially-annotated data: data without action labels and data without text labels. For brevity, we abbreviate them as data w/o action labels and data w/o text labels, respectively, throughout the rest of the paper.
Data w/o action labels contains videos with text annotations but does not include action labels. This type of data can be readily sourced from various public datasets [8, 9, 10]. Data w/o text labels, on the other hand, contains videos with corresponding action labels but does not include text annotations. They can be collected on a large scale by allowing robots to execute policies without constant human supervision [6]. GR-MG is designed to leverage both types of partially-annotated data along with fully-annotated data to maximum the usage of data available.
@@ III-B Network Architecture
The network architecture of GR-MG is illustrated in Fig. 2. It consists of two modules: a progress-guided goal image generation model and a multi-modal goal-conditioned policy.
@@ III-B1 Progress-guided Goal Image Generation Model
This module generates the goal image based on the current observation image, a text description of the task, and the task progress. Specifically, we use InstructPix2Pix [31], a diffusion-based image-editing model, as the network. Following the approach in Susie [14], we generate a sub-goal representing an intermediate state rather than the final state, and update the sub-goal at fixed time intervals during the rollout. However, unlike Susie which generates the goal image solely based on the current observation and the text instruction, we incorporate task progress information in the image generation process. The insight is that robot manipulation is a sequential decision process instead of static image editing. Additionally, the current observation can be ambiguous in the case where there are identical observations throughout the rollout of a task, e.g., moving an object back-and-forth. The progress information can provide valuable global information for goal image generation. During training, the progress of a frame can be easily obtained from its timestep in the trajectory (or video). During inference, such information is not accessible. To address this, we train the policy to predict task progress alongside the action trajectory, as outlined in the following paragraph. To incorporate progress information into the generation model, we append a sentence to the task description, e.g., "pick up the red block. And 60% of the instruction has been completed.". We use T5-Base [32] to encode the text.
III-B2 Multi-modal Goal-Conditioned Policy
@@
Following GR-1 [4], we employ a similar GPT-style transformer network [33]. We briefly review GR-1 for completeness. GR-1 takes as inputs a text instruction, a sequence of observation images, and a sequence of robot states. It outputs actions and future images. The inputs are encoded into tokens with corresponding encoders [34, 35] or linear layers. GR-1 leverages an [ACT] query token and multiple [OBS] query tokens for action and future image prediction, respectively. The query tokens and the input tokens are concatenated together before being fed into the transformer. The output embedding of the [ACT] token are passed through linear layers to predict actions. The output embeddings of the [OBS] tokens are concatenated with learnable mask tokens to predict future images with a transformer. We refer readers to [4] for more details.
@@
GR-MG differs from GR-1 in three ways. Firstly, GR-MG is a multi-modal goal-conditioned policy which takes a goal image and a text instruction as conditions; GR-1 only uses the text for conditioning. To inject the goal image condition, we use MAE [35] to tokenize the goal image generated by the goal image generation model and append the tokens at the front of the input token sequence (Fig. 2). These tokens can be attended by all the subsequent tokens for conditioning. Secondly, in order to predict task progress, a [PROG] query token is included. The output embedding of the token is passed through linear layers to regress the progress value. Finally, we follow recent work [36] to predict action trajectories instead of a single action as in GR-1 via a conditional variational autoencoder (cVAE) [37, 38]. Specifically, we use a VAE encoder to encode the action trajectory into a style vector embedding. We concatenate the style vector embedding, the output embedding of the [ACT] token, and $k$ learnable token embeddings together and input them to a transformer for predicting an action trajectory of $k$ steps.